# Diffusion Dataset Condensation: Training Your Diffusion Model Faster with Less Data

### v2

@@ Abstract
Diffusion models have achieved remarkable success in various generative tasks, but training them remains highly resource-intensive, often with millions of images and GPU days of computation required. From a data-centric perspective addressing the limitation, we study diffusion dataset condensation as a new challenging problem setting that aims at constructing a “synthetic” sub-dataset with significantly fewer samples than the original dataset for training high-quality diffusion models significantly faster. To the best of our knowledge, we are the first to formally study the dataset condensation task for diffusion models, while conventional dataset condensation focused on training discriminative models. For this new challenge, we further propose a novel Diffusion Dataset Condensation (D2C) framework, that consists of two phases: Select and Attach. The Select phase identifies a compact and diverse subset via a diffusion difficulty score and interval sampling, upon which the Attach phase enhances conditional signals and information of the selected subset by attaching rich semantic and visual representations. Extensive experiments across dataset sizes, model architectures, and resolutions demonstrate that our D2C can train diffusion models significantly faster with dramatically fewer data while retaining high visual quality. Notably, for the SiT-XL/2 architecture, our D2C achieves a 100$\times$ acceleration, reaching a FID of 4.3 in just 40k steps using only 0.8% of the training data.
@@
Figure 1: Our D2C framework significantly accelerates diffusion model training with limited data. (a) Overview of our D2C pipeline, which consists of a Select phase that filters a compact and diverse subset via diffusion difficulty score and interval sampling, and an Attach phase that enriches samples with semantic and visual information. (b) D2C achieves over 100$\times$ faster convergence compared to REPA and over 233$\times$ faster than vanilla SiT-XL/2, reaching a FID of 4.3 at just 40k steps. (c) Under a strict 4% data budget (0.05M), our method achieves a FID of 2.7 at 180k iterations, demonstrating its strong training efficiency and rapid convergence.
@@ 1 Introduction
Generative models, such as score-based [1, 2, 3, 4] and flow-based [5, 6, 7] approaches, have achieved remarkable success in various generative tasks, producing high-quality and diverse data across domains [8, 9, 10]. However, these approaches are notoriously data and compute intensive to train, often requiring millions of samples and hundreds of thousands of iterations to capture complex high-dimensional distributions [11, 12, 13]. The resulting cost presents a significant barrier to broader application and iteration within the AIGC community, making efficient training increasingly important across both academic and industrial settings [14, 15]. Recent efforts have improved diffusion training efficiency through various strategies, such as architectural redesigns [13, 16, 11], attention optimization [17], reweighting strategies [18, 19, 20, 8], and representation learning [12]. In parallel, data-centric approaches such as Patch Diffusion [21] and advanced augmentation techniques [8] aim to better exploit the potential of existing data. Despite these advancements, the exploration of building a relatively complete “synthetic” subset through dataset condensation [22] to accelerate diffusion model training with minimal information loss remains underexplored.
@@
Dataset condensation [22, 23, 24, 25] aims to construct a “synthetic” sub-dataset with significantly fewer samples than the original dataset, such that a model trained from scratch on this subset achieves comparable performance to one trained on the full dataset. Unlike data pruning or selection [26], which passively select existing samples, condensation actively optimizes synthetic data, offering greater potential for aggressive data reduction and training efficiency [27]. However, all existing methods are designed for discriminative tasks. Compared to discriminative tasks, generative tasks are much more complex and demand higher dataset quality [28]. Applying popular methods (e.g., SRe2[24]) that have substantiated effective for discriminative tasks to diffusion models presents significant challenges, such as the failure to produce diverse, high-quality outputs with structural and semantic fidelity, leading to degraded results and unstable convergence (see Sec. 4).
@@
We raise a key question: “Can we train diffusion models dramatically faster with significantly less data, while retaining high generation quality?” The answer is affirmative. Answering this question holds significant relevance for the further development of visually generative intelligence and is therefore extremely worthwhile to explore. In this paper, we mainly made the following contributions.
@@
First, to the best of our knowledge, we are the first to formally study the dataset condensation task for diffusion models, a new challenging problem setting that aims at constructing a “synthetic” sub-dataset with significantly fewer samples than the original dataset for training high-quality diffusion models significantly faster. More specifically, our explorations with the diffusion model provide the first insights into the challenges and potential solutions for applying dataset condensation to vision generation tasks. We note that while conventional dataset condensation made great progress and sometimes use diffusion models to construct a subset, this line of research only focused on training discriminative models instead of generative models.
@@
Second, we propose a a novel two-stage Diffusion Dataset Condensation (D2C) framework, consisting of Select and Attach, illustrated in Fig. 1 (a). Our framework addresses the challenges of dataset condensation for diffusion models by decomposing the problem into two key aspects: the Select stage identifies an informative, compact, and learnable subset by ranking samples using the diffusion difficulty score derived from a pre-trained diffusion model; the Attach stage enriches each selected sample by adding semantic and visual representations, further enhancing the training efficiency while preserving performance.
Third, extensive experiments demonstrate that the proposed D2C can train diffusion models significantly faster with dramatically fewer data while retaining high visual quality, substantiating the effectiveness and scalability. Specifically, D2C significantly outperforms random sampling and several popular dataset selection and distillation algorithms (e.g., SRe2[24] and K-Center) across data compression ratios of 0.8%, 4%, and 8%, at resolutions of 256$\times$256 and 512$\times$512, and with both SiT [13] and DiT [11] architectures. In particular, D2C achieves a FID of 4.3 in merely 40k training steps (w/o classifier-free guidance (CFG) [4]) using SiT-XL/2 [13], demonstrating a 100$\times$ acceleration over REPA [12] and a 233$\times$ speed-up compared to vanilla SiT. Furthermore, it further improves to a FID of 2.7 using only 50k synthesized images with CFG (refer to Fig. 1 (c)).
@@ 2 Preliminaries and Related Work
In this section, we briefly review diffusion models as well as dataset condensation.
Diffusion Models. We briefly introduce the standard latent-space noise injection formulation that underlies many diffusion models [6, 11, 9]. These models define a forward process that gradually perturbs input data with Gaussian noise, and train a neural network to approximate the reverse denoising process for sampling (see Appendix B for more details).
Let the original variable be denoted as $\mathbf{x}_{0}\sim q_{0}(\mathbf{x})$. The forward diffusion process is typically implemented as a Markov chain with Gaussian noise injections:
@@
where $\alpha_{t},\ \sigma_{t}\in\mathbb{R}^{+}$ are differentiable functions of $t$ with bounded derivatives. The choice for $\alpha_{t}$ and $\sigma_{t}$ is referred to as the noise schedule of a diffusion model. A commonly used reparameterization in score-based diffusion models [3] expresses the noisy data $\mathbf{x}_{t}$ at time step $t$ as $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\epsilon$, where $\ \epsilon\sim\mathcal{N}(0,\mathbf{I})$. In this case, to recover “clean” data from noise, a neural network $\epsilon_{\theta}(\cdot,\cdot,\cdot)$ is trained to predict the added noise $\epsilon$, thereby approximating the reverse process. This framework adopt the conventional noise prediction strategy for training, where the training objective is to minimize the mean squared error between the predicted and the ground true noise:
@@
Here, $\mathbf{c}$ is a conditional input, such as class labels or text embeddings. In some cases, the prediction target is replaced with a linear combination of noise and data (e.g. $v$-prediction), which corresponds to flow matching [5, 6, 7]:
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{\mathbf{x}_{0}\sim q_{0}(\mathbf{x}),\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\|\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-(\epsilon-\mathbf{x}_{0})\|^{2}_{2}\right].$ (3)
Here, $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ denotes the learned time-dependent velocity field.
Data-centric Efficient Training. Various model-side strategies have been proposed to accelerate diffusion model, including architectural enhancements [17, 11, 13], sampling refinements [18, 19, 20, 8], and representation-level techniques that leverage pretrained vision features [12, 29]. However, relatively little has been explored from a data-centirc perspective. In data-centric model training, given an original dataset $\mathcal{D}=(\hat{\mathbf{X}},\hat{\mathbf{Y}})=\{(\hat{\mathbf{x}}_{i},\hat{y}_{i})\}_{i=1}^{|\mathcal{D}|}$, where each $\hat{y}_{i}$ is the label corresponding to sample $\hat{\mathbf{x}}_{i}$, dataset compression aims to reduce the size of training data while preserving model performance. Two primary strategies have been extensively studied in this context: dataset pruning and dataset condensation.
1) Dataset Pruning. Dataset pruning selects an information-enrichment subset from the original dataset, i.e., $\mathcal{D}^{\text{core}}\subset\mathcal{D}$ with $|\mathcal{D}^{\text{core}}|\ll|\mathcal{D}|$, and directly minimizes the training loss over the subset:
@@
where $\ell(\cdot,\cdot)$ denotes the empirical training loss, and $\phi_{\theta_{\mathcal{D}^{\text{core}}}}$ is the model parameterized by $\theta_{\mathcal{D}^{\text{core}}}$. Classical data pruning methods like random sampling, K-Center [30], and Herding [31] can be used with diffusion models, but they offer minimal performance improvements.
2) Dataset Condensation. In contrast, dataset condensation aims to synthesize a small, compact, and diverse synthetic dataset $\mathcal{D}^{\mathcal{S}}=(\mathbf{X},\mathbf{Y})=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{|\mathcal{D}^{\mathcal{S}}|}$ to replace the original dataset $\mathcal{D}$. The synthetic dataset $\mathcal{D}^{\mathcal{S}}$ is generated by a condensation algorithm $\mathcal{C}$ such that $\mathcal{D}^{\mathcal{S}}\in\mathcal{C}(\mathcal{D})$, with $|\mathcal{D}^{\mathcal{S}}|\ll|\mathcal{D}|$. Each $y_{j}$ corresponds to the synthetic label for the sample $\mathbf{x}_{j}$.
@@
where $\theta_{\mathcal{D}}$ is the parameter set of the neural network $\phi$ optimized on $\mathcal{D}$: $\theta_{\mathcal{D}}=\arg\min_{\theta}\mathbb{E}_{(\hat{\mathbf{x}},\hat{y})\sim\mathcal{D}}\left[\ell(\phi_{\theta}(\hat{\mathbf{x}}),\hat{y})\right].$ A similar definition applies to $\theta_{\mathcal{D}}^{\mathcal{S}}$, which is optimized on the synthetic dataset $\mathcal{D}^{\mathcal{S}}$. Existing methods, such as RDED, MTT, SRe2L, and G-VBSM [27, 24, 25, 32, 33], are primarily designed for discriminative tasks. When applied to diffusion models, these methods generate synthetic images that deviate from the original data distribution, leading to detrimental effects on model training. Visualization of these synthesized images can be found at Appendix I.
Figure 2: Overview of Diffusion Dataset Condensation (D2C). D2C employs a two-stage process: Select and Attach. The Select stage identifies a compact and diverse subset by intervaling sampling using the diffusion difficulty score derived from a pre-trained diffusion model. The Attach stage further enriches each selected sample by adding semantic information and visual information.
@@ 3 Diffusion Dataset Condensation
To enable data-centric efficient training of diffusion models under limited resources, we propose Diffusion Dataset Condensation (D2C), the first unified framework that systematically condenses training data for diffusion models. As illustrated in Fig. 2, this process produces a condensed dataset suitable for efficient diffusion model training. D2C consists of two stages: Select (Sec. 3.1), which identifies a compact set of diverse and learnable real images using diffusion difficulty score and interval sampling techniques; and Attach (Sec. 3.2), which augments each selected image with semantic and visual information to improve generation performance. Finally, we present the novel training paradigm utilizing the condensed dataset generated by D2C (Sec. 3.3).
@@ 3.1 Select: Difficulty-Aware Selection
In this work, we focus on class-to-image (C2I) synthesis. Given a class-conditioned dataset $\mathcal{D}=\bigcup_{y=1}^{C}\mathcal{D}_{y}$, where $\mathcal{D}_{y}=\{x_{i}\}_{i=1}^{|\mathcal{D}_{y}|}$ denotes all samples of class $y$, we aim to select a compact subset for efficient diffusion training. To achieve this, we propose diffusion difficulty score to quantify the denoising difficulty of each sample, followed by our designed interval sampling to ensure diversity within the selected subset.
@@
Diffusion Difficulty Score. The arrangement of samples from easy to hard is crucial for revealing underlying data patterns and facilitating difficulty-aware selection. Recent work [34] demonstrates that diffusion models inherently encode semantic-related class-conditional probability $p_{\theta}(\mathbf{c}|\mathbf{x})$ through the variational lower bound (i.e., diffusion loss Eq. 2) of $\log p_{\theta}(\mathbf{x}|\mathbf{c})$ [3, 1]. This conditional probability can be formulated as $p_{\theta}(\mathbf{c}|\mathbf{x})=\frac{p_{\theta}(\mathbf{x}|\mathbf{c})p(\mathbf{c})}{\sum_{\mathbf{\hat{c}}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})p(\mathbf{\hat{c}})}$. Intuitively, a larger $p_{\theta}(\mathbf{c}|\mathbf{x})$ indicates that sample $\mathbf{x}$ can be more confidently identified as belonging to class $\mathbf{c}$, thus suggesting lower learning difficulty. Given the significant computational overhead of the full Bayesian formulation and our focus on estimating sample difficulty, we ignore the denominator part $\sum_{\mathbf{\hat{c}}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})p(\mathbf{\hat{c}})$ of the calculation. Since the category $y\sim U\{0,\cdots,C\}$ ($C$ denotes the class number) is obtained by uniform sampling, and assuming $\sup\{|\mathbb{E}_{\mathbf{\hat{c}}}[p_{\theta}(\mathbf{x_{1}}|\mathbf{\hat{c}})]-|\mathbb{E}_{\mathbf{\hat{c}}}[p_{\theta}(\mathbf{x_{2}}|\mathbf{\hat{c}})]|\}_{\mathbf{x_{1}},\mathbf{x}_{2}\sim\mathcal{D^{X}}}\leq\eta$ ($\mathcal{D^{X}}$ denotes the all images in the dataset), we define the diffusion difficulty score based on the class-conditional probability $p_{\theta}(\mathbf{x}|\mathbf{c})\propto\frac{p_{\theta}(\mathbf{x}|\mathbf{c})}{\sum_{\mathbf{c}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})}=p_{\theta}(\mathbf{c}|\mathbf{x})$ ($\sum_{\mathbf{c}}p_{\theta}(\mathbf{x}|\mathbf{\hat{c}})$ can be viewed as a constant when sorting by sample difficulty):
$s_{\text{diff}}(\mathbf{x})=-p_{\theta}(\mathbf{x}|\mathbf{c})=-\log\left(\exp\left(-\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[\left\|\epsilon-\epsilon_{\theta}\left(\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon,\,t,\,\mathbf{c}\right)\right\|^{2}_{2}\right]\right)\right),$ (6)
@@
The higher the score $s_{\text{diff}}(\mathbf{x})$, the easier it is, and the lower the score $s_{\text{diff}}(\mathbf{x})$, the more difficult it is. Note that in this paper’s figures and tables, we directly use the diffusion loss (i.e., $\propto-p_{\theta}(\mathbf{c}|\mathbf{x})$) for ease of understanding.
By computing $s_{\text{diff}}(x)$ for all training samples, we construct a ranked dataset. As shown in Fig. 3 (Left), these scores exhibit a skewed unimodal distribution. Selecting the easiest samples (Min) yields a subset dominated by clean, background-simple images with high learnability but limited diversity. In contrast, selecting only the highest-score samples (Max) results in cluttered, noisy, and ambiguous images that are difficult to optimize. Meanwhile, many samples lie in the middle range, offering moderate learnability but richer contextual information.
Interval Sampling. To balance diversity and learnability, we propose an interval sampling strategy. Specifically, we sort its images $\mathcal{D}_{y}$ within each class $y$ in ascending order of $s_{\text{diff}}(x)$ and select samples at a fixed interval $k$:
$\mathcal{D}_{\text{IS}}=\bigcup_{y=1}^{C}\left\{x^{(i)}\in\mathcal{D}_{y}\;\middle|\;i\in\{0,k,2k,\dots\}\right\},$ (7)
where $\mathcal{D}_{\text{IS}}$ denotes the selected subset constructed by interval sampling, $k$ is the fixed sampling interval, and $x^{(i)}$ is the $i$-th sample in the sorted list (e.g., $x^{(0)}$ corresponds to the sample with the lowest diffusion difficulty score). Interval sampling with a larger interval $k$ promotes diversity in the sampled data while potentially hindering learnability. As shown in Fig. 3 (Left), this trade-off arises from a shift in the sample distribution: a larger $k$ leads to a reduction in the number of easy samples and a corresponding increase in the representation of standard and difficult samples.
@@
Following interval sampling yields notable findings. First, interval sampling efficiently mitigates the inclusion of both overly simplistic and excessively challenging samples, yielding subsets that exhibit both learnability and diversity. Fig. 3 (right) displays representative examples selected by our strategy, showcasing structural clarity coupled with contextual richness. Second, the FID of diffusion models trained on the subset $\mathcal{D}_{\text{IS}}$ generated with increasing $k$ initially decreases and subsequently increases. For example, as shown in Fig. 8, our results further corroborate that $k=16$ strikes a favorable balance between sample diversity and generation quality when using a 50K data budget.

### v3

# Accelerating Diffusion Model Training under Minimal Budgets: A Condensation-Based Perspective

@@ Abstract
[…]
@@
[…]
@@ 1 Introduction
Generative models, such as score-based [43, 42, 17] and flow-based [27] approaches, have achieved remarkable success in various generative tasks [3], producing high-quality and diverse data across domains [19, 12, 4]. However, these approaches are notoriously data and compute intensive to train, often requiring millions of samples and hundreds of thousands of iterations to capture complex high-dimensional distributions [33, 53, 28]. The resulting cost presents a significant barrier to broader application and iteration within the AIGC community, making efficient training increasingly important across both academic and industrial settings [41, 26, 11]. Recent efforts have improved diffusion training efficiency through various strategies, such as architectural redesigns [28, 55, 33], attention optimization [1, 24], reweighting strategies [13], and representation learning [53, 48, 54]. In parallel, data-centric approaches such as patch-based methods [7, 47], Infobatch [34] and Reweighting[25] aim to better exploit the potential of existing data. Despite these advances, directly constructing a much smaller yet informative training subset as a primary condensation-based route to accelerate diffusion training remains largely unexplored, even though it provides a particularly direct and effective way to reduce data budgets.
@@
Dataset condensation [49, 46, 52, 38] aims to construct a much smaller condensed sub-dataset with significantly fewer samples than the original dataset, such that a model trained from scratch on this subset achieves performance comparable to one trained on the full dataset while converging much faster. In practice, existing DC methods typically instantiate this objective in two ways [49]: (i) pixel-level dataset distillation, which directly optimizes synthetic images [52, 2], and (ii) image-level condensation, which operates on real images through selection and transformation [44, 10, 20]. Unlike classical data pruning or selection [51], which passively select a fixed subset of existing samples, both branches of dataset condensation actively construct condensed training sets, either by optimizing synthetic images or by selecting and enriching real ones, thereby enabling more aggressive data reduction and higher training efficiency [44]. However, these methods have been developed almost exclusively for discriminative tasks. Compared to discriminative learning, generative diffusion training is substantially more complex and demands higher dataset quality [29]; directly applying popular DC algorithms (e.g., SRe2[52], RDED [44]) to diffusion models often leads to synthetic images that lack structural and semantic fidelity, resulting in degraded sample quality and unstable convergence (see Sec. 4).
@@
We raise a key question: “Can we train diffusion models dramatically faster with significantly less data, while retaining high generation quality?” The answer is affirmative. In this paper, we make three main contributions.
@@
First, to the best of our knowledge, we are the first to formally study the dataset condensation task for diffusion models, a new challenging problem setting that aims at constructing a “condensed” sub-dataset with significantly fewer samples than the original dataset for training high-quality diffusion models significantly faster. We address a fundamental academic gap concerning the application of dataset condensation in diffusion models. More specifically, our explorations with the diffusion model provide the first insights into the challenges and potential solutions for applying dataset condensation to vision generation tasks. We note that while conventional dataset condensation has made great progress and sometimes uses diffusion models to construct a subset, this line of research only focused on training discriminative models instead of generative models.
@@
Second, we propose D2C, a novel two-stage dataset condensation framework tailored for training diffusion models. Our framework addresses the challenges of dataset condensation for diffusion models by decomposing the problem into two key aspects: the Select stage identifies an informative, compact, and learnable subset by ranking samples using the diffusion difficulty score derived from a pre-trained diffusion model; the Attach stage enriches each selected sample by adding semantic and visual representations, further enhancing the training efficiency while preserving performance.
[…]
@@ 2 Preliminaries and Related Work
Diffusion Models. We briefly introduce the standard latent-space noise injection formulation [33], which defines a forward process that gradually perturbs input data $\mathbf{x}_{0}\sim q_{0}(\mathbf{x})$ with Gaussian noise:
@@
where $\alpha_{t},\ \sigma_{t}\in\mathbb{R}^{+}$ are differentiable functions of $t$ with bounded derivatives. The choice for $\alpha_{t}$ and $\sigma_{t}$ is referred to as the noise schedule of a diffusion model. After that, we need to train a neural network $\epsilon_{\theta}(\cdot,\cdot,\cdot)$ to approximate the reverse denoising process (i.e., predict the added noise $\epsilon$) for sampling (see Appendix B for more details). The training objective is to minimize the mean squared error between the predicted and the ground true noise:
@@
Here, $\mathbf{c}$ is a conditional input, such as class labels or text embeddings. In some cases, the prediction target is replaced with the $v$-prediction, which corresponds to flow matching.
Figure 2: Overview of Diffusion Dataset Condensation (D2C). D2C employs a two-stage process: Select and Attach. The Select stage identifies a compact and diverse subset by interval sampling using the diffusion difficulty score derived from a pre-trained diffusion model. The Attach stage further enriches each selected sample by adding semantic information and visual information.
Data-centric Efficient Training. Beyond model-side improvements, a complementary line of work takes a data-centric view on diffusion efficiency. Patch-based schemes [7, 47] and Infobatch [34] focus on reallocating training effort over existing samples by reweighting or resampling informative regions and instances. However, comparatively few methods directly tackle diffusion training efficiency by explicitly reducing and restructuring the overall training set. In this setting, given an original dataset $\mathcal{D}=\{(\hat{\mathbf{x}}_{i},\hat{y}_{i})\}_{i=1}^{|\mathcal{D}|}$, where each $\hat{y}_{i}$ is the label corresponding to sample $\hat{\mathbf{x}}_{i}$, dataset compression aims to reduce the size of training data while preserving model performance. Two primary strategies have been extensively studied in this context: dataset pruning and dataset condensation.
1) Dataset Pruning. Dataset pruning selects an information-enriched subset from the original dataset, i.e., $\mathcal{D}^{\text{core}}\subset\mathcal{D}$ with $|\mathcal{D}^{\text{core}}|\ll|\mathcal{D}|$, and directly minimizes the training loss over the subset:
@@
where $\ell(\cdot,\cdot)$ denotes the empirical training loss, and $\phi_{\theta_{\mathcal{D}^{\text{core}}}}$ is the model parameterized by $\theta_{\mathcal{D}^{\text{core}}}$. Classical data pruning methods like random sampling, K-Center [18], and Herding [5] can be used with diffusion models, but they offer minimal performance improvements. Very recently, Li et al. [25] investigate data-efficient diffusion training from the perspective of dataset pruning by selecting a coreset with surrogate features and then performing class-wise reweighting. While this approach substantially reduces training cost and improves over naive pruning, it does not attach any additional information to the selected samples and is mainly validated on relatively small-scale or latent diffusion settings, which limits its ability to fully exploit the potential of condensed training data for large-scale, high-resolution diffusion models.
2) Dataset Condensation. Following recent work [49], dataset condensation aims to synthesize a small, compact, and diverse synthetic dataset $\mathcal{D}^{\mathcal{S}}=(\mathbf{X},\mathbf{Y})=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{|\mathcal{D}^{\mathcal{S}}|}$ to replace the original dataset $\mathcal{D}$. The synthetic dataset $\mathcal{D}^{\mathcal{S}}$ is generated by a condensation algorithm $\mathcal{C}$ such that $\mathcal{D}^{\mathcal{S}}\in\mathcal{C}(\mathcal{D})$, with $|\mathcal{D}^{\mathcal{S}}|\ll|\mathcal{D}|$. Each $y_{j}$ corresponds to the synthetic label for the sample $\mathbf{x}_{j}$.
@@
[…]
@@ 3 Diffusion Dataset Condensation
As illustrated in Fig. 2, D2C consists of two stages: Select (Sec. 3.1), which identifies a compact set of diverse and learnable real images using diffusion difficulty score and interval sampling techniques; and Attach (Sec. 3.2), which augments each selected image with semantic and visual information to improve generation performance. Finally, we describe how to train diffusion models on the condensed dataset produced by D2C in Sec. 3.3. (Sec. 3.3).
@@ 3.1 Select: Difficulty-Aware Selection
In this work, we focus on class-to-image (C2I) synthesis, aligned with the setting in Yu et al. [53], and show that our framework also applies to the text-to-image (T2I) setting with only minor changes; see Appendix G for details. Given a class-conditioned dataset $\mathcal{D}=\bigcup_{y=1}^{C}\mathcal{D}_{y}$, where $C$ denotes the class number and $\mathcal{D}_{y}=\{x_{i}\}_{i=1}^{|\mathcal{D}_{y}|}$ denotes all samples of class $y$, our aim is to select a compact subset for efficient diffusion training. To achieve this, we propose the diffusion difficulty score to quantify the denoising difficulty of each sample, followed by our designed interval sampling to ensure diversity within the selected subset.
@@
Diffusion Difficulty Score. The arrangement of samples from easy to hard is crucial for revealing underlying data patterns and facilitating difficulty-aware selection. Recent work [23, 56] demonstrates that diffusion models inherently encode semantic-related class-conditional probability $p_{\theta}(\mathbf{c}\mid\mathbf{x})$ through the variational lower bound (i.e., diffusion loss Eq. 2) of $\log p_{\theta}(\mathbf{x}\mid\mathbf{c})$ [17, 43]. This conditional probability admits the standard Bayesian form
$p_{\theta}(\mathbf{c}\mid\mathbf{x})=\frac{p_{\theta}(\mathbf{x}\mid\mathbf{c})\,p(\mathbf{c})}{\sum_{\hat{c}}p_{\theta}(\mathbf{x}\mid\hat{c})\,p(\hat{c})}$. (5)
Intuitively, a larger $p_{\theta}(\mathbf{c}\mid\mathbf{x})$ indicates that sample $\mathbf{x}$ can be more confidently identified as belonging to class $\mathbf{c}$, thus suggesting lower learning difficulty. Computing the full denominator in Eq. (5) for every sample is expensive, while we only need a score that orders samples by difficulty. Since the class label $y\sim U\{1,\dots,C\}$ is obtained by uniform sampling and the average likelihood over classes does not vary too much across samples, i.e., we assume $\sup_{\mathbf{x}_{1},\mathbf{x}_{2}\sim\mathcal{D}}\bigl|\mathbb{E}_{\hat{c}}[p_{\theta}(\mathbf{x}_{1}\mid\hat{c})]-\mathbb{E}_{\hat{c}}[p_{\theta}(\mathbf{x}_{2}\mid\hat{c})]\bigr|\leq\eta$,where $\eta>0$ is a small tolerance, the denominator in Eq. (5) can be treated as approximately constant with respect to $\mathbf{x}$. Consequently, the posterior is proportional to the class-conditional likelihood,
$p_{\theta}(\mathbf{c}\mid\mathbf{x})\propto p_{\theta}(\mathbf{x}\mid\mathbf{c}).$ (6)
We define the diffusion difficulty score based on this posterior:
$s_{\text{diff}}(\mathbf{x})$ $=-p_{\theta}(\mathbf{c}\mid\mathbf{x})\propto-p_{\theta}(\mathbf{x}\mid\mathbf{c})$ (7)
$=-\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I}),\,t\sim\mathcal{U}[0,1]}\left[\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{x_{t}},t,\mathbf{c}\right)\right\|^{2}_{2}\right]$
@@
The higher the score $s_{\text{diff}}(\mathbf{x})$, the more difficult it is, and the lower the score $s_{\text{diff}}(\mathbf{x})$, the easier it is. To simplify our presentation, we define the diffusion loss $p_{\theta}(\mathbf{x}|\mathbf{c})$ as the diffusion difficulty score.
@@
By computing $s_{\text{diff}}(x)$ for all training samples, we construct a ranked dataset. As shown in Fig. 3, these scores exhibit a skewed unimodal distribution. Selecting the easiest samples (Min) yields a subset dominated by clean, background-simple images with high learnability but limited diversity. In contrast, selecting only the highest-score samples (Max) results in cluttered, noisy, and ambiguous images that are difficult to optimize. Meanwhile, many samples lie in the middle range, offering moderate learnability but richer contextual information. Selecting an appropriate value within this range is therefore critical; we provide a more detailed discussion in Appendix H.1.
Interval Sampling. To balance diversity and learnability, we propose an interval sampling strategy. Specifically, we sort its images $\mathcal{D}_{y}$ within each class $y$ in ascending order of $s_{\text{diff}}(x)$ and select samples at a fixed interval $k$: $\mathcal{D}_{\text{IS}}=\bigcup_{y=1}^{C}\left\{x^{(i)}\in\mathcal{D}_{y}\;\middle|\;i\in\{0,k,2k,\dots\}\right\}$, where $\mathcal{D}_{\text{IS}}$ denotes the selected subset constructed by interval sampling, $k$ is the fixed sampling interval, and $x^{(i)}$ is the $i$-th sample in the sorted list (e.g., $x^{(0)}$ corresponds to the sample with the lowest diffusion difficulty score). Interval sampling with a larger interval $k$ promotes diversity in the sampled data while potentially hindering learnability. As shown in Fig. 3 (Left), this trade-off arises from a shift in the sample distribution: a larger $k$ leads to a reduction in the number of easy samples and a corresponding increase in the representation of standard and difficult samples.
Figure 4: Overview of DC-Embedding.
Extended Discussion. Training exclusively on the easiest (Min) or the hardest (Max) samples is suboptimal. Instead, a balanced curriculum comprising easy, medium, and difficult examples yields a training subset that is both learnable and diverse, ultimately leading to stronger generative performance. We further offer more discussions and insights on interval sampling in Appendix H.2.